OpenStudy (baldymcgee6):

Determine the following features of the line that contains the points (-1, 1) and (2, 5): a) Its slope b) Its point-slope equation c) Its slope-intercept equation d) Its general equation

OpenStudy (anonymous):

first find slope m=slope=change in y values/ change in x values m=(5-1)/(2-(-1)=4/3 point slope form is y-y0=m(x-x0) y-1=4/3(x-(-1) y-1=4/3(x+1) y-1=4/3x+4/3 (point slope form) simplify it for y to get slope intercept form. y=4/3x+4/3+1 y=4/3x+7/3 (slope intercept form) 4/3x-y+7/3=0 multiply with 3 whole equation. 4x-3y+7=0 (general form.)

OpenStudy (baldymcgee6):

thank youuuu